OpenStudy (anonymous):

Multiply. State any excluded values. 5/2a*3/a^2

OpenStudy (anonymous):

5/2a*3/a^2 \[\frac{ 5*3 }{ 2a*a^{2} } = \frac{ 15 }{ 2a ^{3} }\]

OpenStudy (anonymous):

ty maybe you can help on my next one

OpenStudy (anonymous):

sure